A) Determine the resistance of a 75-W lightbulb connected to its proper source voltage of 110 V.


B) Determine the current through a 75-W lightbulb connected to its proper source voltage of 110 V.


C) Determine the resistance of a 250-W lightbulb connected to its proper source voltage of 110 V.


D) Determine the current through a 250-W lightbulb connected to its proper source voltage of 110 V.         
Answer #1

b) I = 75/110 = 0.682 amps



a) R = 110/0.682 = 161 ohms



d) I=250/110 = 2.27 amps


c) R = 110/2.27 = 48.45 ohms
Answer #2

v2 1102 P 75 A) Resistance of the bulb is given by,R- 161.332 B) Current through the bulb is, I = -0.68A VR V161.33 v2 110* P
Answer #3

A) P= V^2/R so, R= V^2/P , therefore, R= 110*110/75 = 161.33 ohm


B) P= VI so I =P/V , therefore , I =75/110 = .68 Ampere


C) P= V^2/R so, R= V^2/P , therefore, R= 110*110/250 = 48.4 ohm


B) P= VI so I =P/V , therefore , I =250/110 = 2.272 Ampere